Minmax Co.'s direct labor information for February is as follows: Direct labor hours worked (AQ) 37,000 Standard direct labor hours for units manufactured (SQ) 38,000 Unfavorable direct labor rate variance $ 11,100 Total payroll for direct labor $ 547,600 The total direct labor flexible-budget variance in February, to the nearest dollar, was:
Answer #1
Actual direct labor rate = Total payroll for direct labor / Direct labor hours worked (AQ) = 547600 / 37000 14.80
Unfavorable variance is written as positive value.
Unfavorable direct labor rate variance = ( Actual rate - Standard rate ) * Actual hours
11100 = ( 14.80 - Standard rate ) * 37000
11100 / 37000 = 14.80 - Standard rate
0.30 = 14.80 - Standard rate
Standard rate = 14.80 - 0.30 14.50
Total direct labor flexible budget variance = ( Actual rate * Actual hours ) - ( Standard rate * Standard hours ) = ( 14.80 * 37000 ) - ( 14.50 * 38000 ) 3400 F